# Janus-faces of Temporal Constraint Languages: A dichotomy of expressivity

### v1

@@ 3.1. Templates with pppp
Let $E$ be a binary smooth relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$, and assume that $E$ contains a min-clean tuple $\binom{u^{1}}{v^{1}}$. Then there exist $I\subseteq k$, $m,n\geq 1,q\in\mathbb{Q}$, and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in E$ such that for all $u,v\in\mathbb{Q}^{k}$ the tuple defined by
@@
If $\min u^{1}=\min v^{1}$ then $\operatorname{\textnormal{minx}}u^{1}=\operatorname{\textnormal{minx}}v^{1}$. In this case, we set $I\coloneqq\operatorname{\textnormal{minx}}u^{1}$, $m\coloneqq 1$, $n\coloneqq 1$, and $q\coloneqq\min u^{1}$.
Otherwise, assume without loss of generality that $\min u^{1}<\min v^{1}$. We set $q\coloneqq\min u^{1}$, and inductively set $v^{i+1}\coloneqq u^{i}$, and pick $u^{i+1}$ arbitrarily so that $\binom{u^{i+1}}{v^{i+1}}\in T$. Set $J^{1}\coloneqq\operatorname{\textnormal{minx}}u^{1}$, and inductively $J^{i+1}\coloneqq\{j\;|\;u^{i+1}_{j}\leq q\}\setminus J^{1}\cup\cdots\cup J^{i}$. Finally, we pick $n\geq 1$ minimal such that $J^{n}$ is empty, set $I$ to be the union $J^{1}\cup\cdots\cup J^{n}$, and $m\coloneqq|I|$.
For all $i$ we have $u^{\prime}\sim_{J_{i}}u^{i}$ and $v^{\prime}\sim_{J_{i}}v^{i+1}=u^{i}$, hence $u^{\prime}\sim_{J_{i}}v^{\prime}$. Moreover $\textnormal{pr}_{J_{i}}(u^{\prime})<\textnormal{pr}_{J_{i+1}}(u^{\prime})$, and similarly $\textnormal{pr}_{J_{i}}(v^{\prime})<\textnormal{pr}_{J_{i+1}}(v^{\prime})$ for all $i$. It follows that $u^{\prime}\sim_{I}v^{\prime}$. Moreover, $\textnormal{pr}_{I}(u^{\prime})<\textnormal{pr}_{[k]\setminus I}(u^{\prime})$ and $\textnormal{pr}_{I}(v^{\prime})<\textnormal{pr}_{[k]\setminus I}(v^{\prime})$, hence $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$. Finally, it is easy to verify that $u^{\prime}\sim_{[k]\setminus I}u$ and $v^{\prime}\sim_{[k]\setminus I}v$. ∎
A similar statement holds true for higher-arity relations $R$ that are symmetric. In fact, provided the existence of min-clean tuples in $R$, it suffices that $R$ is cyclic.
@@
Let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$, and assume that $R$ contains a min-clean tuple $u^{1}=(u^{1}_{1}\dots,u^{1}_{n})$. Then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u=(u_{1},\dots,u_{n})\in R$ the tuple $u^{\prime}=(u^{\prime}_{1},\dots,u^{\prime}_{n})$ defined by
@@
Without loss of generality assume that $1\in M(u^{1})$, and set $q\coloneqq\min(u^{1}_{1})$. By min-cleanliness of $u^{1}$, its cyclic permutations $u^{2}\coloneqq(u^{1}_{n},u^{1}_{1},u^{1}_{2},\dots),\dots,u^{n}\coloneqq(u^{1}_{2},\dots,u^{1}_{1})$ and $I\coloneqq\operatorname{\textnormal{minx}}(u^{1}_{1})$ now have the required properties. ∎
We now get a pseudo-loop by inductively shrinking $k$. To ensure the existence of min-clean tuples in cyclic relations $R$ of arity $n\geq 3$, as shown in Section 5, we additionally require that $R$ be $\Sigma$-invariant for some $2$-transitive subgroup $\Sigma$ of $\textnormal{Sym}(n)$.
@@
We use induction on $k$. The claim is trivial for $k=1$. Assume we have already shown it for all $j<k$. Lemmata 5.1, 5.5, 5.9, 5.2, 5.7, and 5.10 provide min-clean tuples in all cases, and the fact that $pp$ is contained in all three clones in question allows us to use Lemmata 3.3 and 3.4, respectively. We prove the statement for the binary case; the case $n\geq 3$ follows by analogy. Let $\binom{u^{1}}{v^{1}}\in R$ be min-clean, and take $I,m,n,q,$ and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in R$ as in Lemma 3.3. By induction hypothesis applied to the relation $\textnormal{pr}_{[k]\setminus I}R$, there exists $\binom{u}{v}\in R$ such that $u\sim_{[k]\setminus I}v$. For $\binom{u^{\prime}}{v^{\prime}}$ defined as in Lemma 3.3, we then have:
@@ 3.2. Templates with $\ell\ell$
In order to adapt Lemma 3.3 to the setting of digraphs preserved by $\ell\ell$, we need to restrict to tuples that are ‘as injective as possible’.
@@
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. For $S\subseteq R$, we set
$S^{\prime}\coloneqq\{t\in\langle S\rangle_{\operatorname{\textnormal{lex}}}\;|\;\ker t=\bigcap_{s\in S}\ker s\}.$
The relation $S^{\prime}$ inherits desirable properties from $S$:
@@
If $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{\textnormal{lex}}\rangle$. If $S\subseteq R$ is non-empty, then $S^{\prime}\neq\varnothing$. If $n=2$ and $S$ is smooth, then so is $S^{\prime}$. If $\Sigma$ is a subgroup of $\textnormal{Sym}(n)$ and $S$ is $\Sigma$-invariant, then so is $S^{\prime}$.
@@
The first statement is obtained by applying $\operatorname{\textnormal{lex}}^{[m]}$} of a sufficiently high arity $m$ to tuples in $S$ until the kernel of the resulting tuple can no more be refined. Suppose that $n=2$ and $S$ is smooth. To see that $S^{\prime}$ is smooth, pick $t\in S^{\prime}$, $f\in\langle\operatorname{\textnormal{lex}}\rangle$ of some arity $m\in\mathbb{N}$, and $s^{1},\dots,s^{m}\in S$ such that $t=f(s^{1},\dots,s^{m}).$ We show that $t_{1}\in\textnormal{pr}_{2}(S^{\prime})$, the converse statement for $t_{2}$ is shown analogously. By smoothness of $S$, for every $i\leq m$ there exists $r^{i}\in\mathbb{Q}^{k}$ such that $\binom{r^{i}}{s^{i}_{1}}\in S$. Let $w\coloneqq f(r^{1},\dots,r^{m})$. Picking any $\binom{u}{v}\in S^{\prime}$, we now have that $\operatorname{\textnormal{lex}}(\binom{w}{t_{1}},\binom{u}{v})\in S^{\prime}$. From $t_{1}\sim_{k}\operatorname{\textnormal{lex}}(t_{1},v)$, the statement follows. The last statement is clear from the definition of $S^{\prime}$. ∎
The following is an adaptation of Lemma 3.3 to the setting of smooth binary relations preserved by $\langle\ell\ell\rangle$.
@@
Again, the existence of min-clean tuples as provided by Lemma 5.13 directly implies the existence of a pseudo-loop.
@@
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$. If $S\subseteq E$ is smooth, has pseudo-algebraic length $1$, and $\textnormal{pr}_{1}(S)$ is weakly connected, then $S^{\prime}$ contains a pseudo-loop.
@@
We proceed with the corresponding statement for hypergraphs preserved by $\langle\ell\ell\rangle$ that are $\Sigma$-invariant for some $2$-transitive cyclic group $\Sigma$.
@@
Once more, we proceed by induction on $k$, employing Lemma 5.15 for the existence of a min-clean tuple within $R^{\prime}$, and the slight adaption of Lemma 3.4 to the setting of $\ell\ell$ changing $R$ to $R^{\prime}$. ∎
3.3. Proof of the main theorems
@@
Replacing $\mathbb{G}$ by the preimage $\bar{\mathbb{G}}$ of $\mathbb{G}$ under the map of the pp-interpretation in $\mathbb{A}$, we may assume that $\mathbb{G}$ is a digraph defined on a finite power of $\mathbb{Q}$ whose edge relation $E$ is preserved by every polymorphism of $\mathbb{A}$. Indeed, by definition of a pp-interpretation, the preimage $\bar{E}$ of $E$ is pp-definable in $\mathbb{A}$, and therefore preserved by $\operatorname{\textnormal{Pol}}(\mathbb{A})$ [BN06]. A pseudo-loop in $\bar{E}$ yields a pseudo-loop in $E$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$. Moreover, $\bar{E}$ is smooth and has pseudo-algebraic length modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$.
By Theorem 2.3, $\mathbb{A}$ is preserved by one of $\ell\ell$, $\min$, $\operatorname{mi}$, $\operatorname{mx}$, their duals, or a constant operation. Accordingly, $\mathbb{G}$ is preserved by one of these operations, and thus by all operations in its generated clone. Clearly, every non-empty relation preserved by a constant operation contains even a loop. Moreover, $\mathbb{G}$ contains a pseudo-loop if and only if $\mathbb{G}$ does. Recalling that, for $f\in\{\ell\ell,\min,\operatorname{mi},\operatorname{mx}\}$, $\mathbb{G}$ is preserved by $\langle\operatorname{dual}(f)\rangle$ if and only if $\mathbb{G}$ is preserved by $\langle f\rangle$, this allows us to restrict ourselves to the cases where $\mathbb{G}$ is preserved by one of $\langle\ell\ell\rangle$, $\langle\min\rangle$, $\langle\operatorname{mi}\rangle$, and $\langle\operatorname{mx}\rangle$. In the first case, let $W$ be a weakly connected component of $\mathbb{G}$ of pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$, and set $S\coloneqq E|_{W\times W}$. By Lemma 3.9, $S^{\prime}\subseteq E$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$. In all the other cases, we can directly apply Lemma 3.5, yielding a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ within $E$. ∎
Similarly, putting together Lemmata 3.10 and 3.5, we obtain the corresponding pseudo-loop lemma for symmetric hypergraphs $\mathbb{T}$. By what we have shown, it in fact suffices for $\mathbb{T}$ to be cyclic and $2$-transitive. This amounts to a strictly weaker condition than full symmetry: for example, for every odd number $n\geq 7$, the alternating group $A_{n}$ is a proper subgroup of $\textnormal{Sym}(n)$ with these properties. It might be of interest to point out that pseudo-loop lemmata involving symmetry conditions provided by non-trivial group actions have also been established for finite hypergraphs (see [BK12, Bru23]). See 1.6
@@ 4. Pseudo-loop conditions from pseudo-loops
Following the standard method – established first by Siggers [Sig10] for finite structures, and adapted to $\omega$-categorical ones by Barto-Pinsker [BP20] – the existence of pseudo-loops in Theorems 1.3 and 1.6 implies the validity of the corresponding pseudo-loop conditions in all temporal constraint languages that are not omni-expressive. See 1.4
@@
Suppose that $\mathbb{G}=(\{1,\dots,n\};E_{\mathbb{G}})$ is a finite smooth digraph of algebraic length $1$, and $u\circ s(x_{i_{1}},\dots,x_{i_{m}})\approx v\circ s(x_{j_{1}},\dots,x_{j_{m}})$ is the corresponding $\mathbb{G}$-pseudo-loop condition. We show that for every $k\geq 1$ and for every $a_{1},\dots,a_{n}\in\mathbb{Q}^{k}$ there exists an $m$-ary operation $s\in\operatorname{\textnormal{Pol}}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit. A standard compactness argument (as provided, for example, in [BP20, Lemma 4.2]) using oligomorphicity of $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ then yields the ‘global’ validity in $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of the desired pseudo-identity.
To this end, take $k\geq 1$, and let $a_{1},\dots,a_{n}$ be arbitrary $k$-tuples of elements of $\mathbb{Q}$. Consider the binary relation $R\subseteq\mathbb{Q}^{k}\times\mathbb{Q}^{k}$ that consists precisely of all the tuples $(a_{i},a_{j})$ for which $(i,j)\in E_{\mathbb{G}}$, and let $E_{\mathbb{H}}\coloneqq\langle R\rangle_{\operatorname{\textnormal{Pol}}(\mathbb{A})}$. Since $E_{\mathbb{H}}$ is, by construction, preserved by all polymorphisms of $\mathbb{A}$, it is pp-definable in $\mathbb{A}$ [BN06]. It follows that $\mathbb{H}\coloneqq(A^{k};E_{\mathbb{H}})$ is pp-interpretable in $\mathbb{A}$. Moreover, observe that $\mathbb{H}$ is a smooth digraph of algebraic length $1$. Applying Theorem 1.3, we get a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ in $\mathbb{H}$. Unravelling definitions, this means that there exists $s\in\operatorname{\textnormal{Pol}}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit. ∎
The corresponding statement for loop conditions induced by symmetric hypergraphs is derived from Theorem 1.6 in a similar way. Again, we may replace the word ‘symmetric’ by ‘cyclic and $2$-transitive’. See 1.7

### v2

@@ 3.1. Templates with pppp
Let $E$ be a binary smooth relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$. If $E$ contains a min-clean tuple $\binom{u^{1}}{v^{1}}$, then there exist $I\subseteq k$, $m,n\geq 1,q\in\mathbb{Q}$, and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in E$ such that for all $u,v\in\mathbb{Q}^{k}$ the tuple defined by
@@
If $\min u^{1}=\min v^{1}$ then $\operatorname{minx}u^{1}=\operatorname{minx}v^{1}$. In this case, we set $I\coloneqq\operatorname{minx}u^{1}$, $m\coloneqq 1$, $n\coloneqq 1$, and $q\coloneqq\min u^{1}$.
Otherwise, assume without loss of generality that $\min u^{1}<\min v^{1}$. We set $q\coloneqq\min u^{1}$, and inductively set $v^{i+1}\coloneqq u^{i}$, and pick $u^{i+1}$ arbitrarily so that $\binom{u^{i+1}}{v^{i+1}}\in T$. Set $J^{1}\coloneqq\operatorname{minx}u^{1}$, and inductively $J^{i+1}\coloneqq\{j\;|\;u^{i+1}_{j}\leq q\}\setminus J^{1}\cup\cdots\cup J^{i}$. Finally, we pick $n\geq 2$ minimal such that $J^{n}$ is empty, set $I$ to be the union $J^{1}\cup\cdots\cup J^{n}$, and $m\coloneqq|I|$.
For all $i$ we have $u^{\prime}\sim_{J_{i}}u^{i}$ and $v^{\prime}\sim_{J_{i}}v^{i+1}=u^{i}$, hence $u^{\prime}\sim_{J_{i}}v^{\prime}$. Moreover $\operatorname{pr}_{J_{i}}(u^{\prime})<\operatorname{pr}_{J_{i+1}}(u^{\prime})$, and similarly $\operatorname{pr}_{J_{i}}(v^{\prime})<\operatorname{pr}_{J_{i+1}}(v^{\prime})$ for all $i$. It follows that $u^{\prime}\sim_{I}v^{\prime}$. Moreover, $\operatorname{pr}_{I}(u^{\prime})<\operatorname{pr}_{[k]\setminus I}(u^{\prime})$ and $\operatorname{pr}_{I}(v^{\prime})<\operatorname{pr}_{[k]\setminus I}(v^{\prime})$, hence $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$. Finally, it is easy to verify that $u^{\prime}\sim_{[k]\setminus I}u$ and $v^{\prime}\sim_{[k]\setminus I}v$. ∎
A similar statement holds true for higher-arity relations $R$ that are cyclic:
@@
Let $n\geq 2$, and let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$. If $R$ contains a min-clean tuple $u^{1}$, then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u\in R$ the tuple $u^{\prime}$ defined by
@@
Without loss of generality assume that $1\in M(u^{1})$, and set $q\coloneqq\min(u^{1}_{1})$. By min-cleanliness of $u^{1}$, its cyclic permutations $u^{2}\coloneqq(u^{1}_{2},u^{1}_{3},\dots,u^{1}_{1}),\dots,u^{n}\coloneqq(u^{1}_{n},u^{1}_{1},u^{1}_{2},\dots)$ and $I\coloneqq\operatorname{minx}(u^{1}_{1})$ now have the required properties. ∎
In both cases, we now get a pseudo-loop by inductively shrinking $k$. To ensure the existence of min-clean tuples in cyclic relations $R$ of arity $n\geq 3$, as shown in Section 6, we need the additional requirement of 2-transitivity.
@@
We use induction on $k$. The claim is trivial for $k=1$. Assume we have already shown it for all $j<k$. Lemmata 6.6, 6.3, 6.9, 6.7, 6.5, and 6.10 provide min-clean tuples in all cases, and the fact that $pp$ is contained in all three clones in question allows us to use Lemmata 3.3 and 3.4, respectively. We prove the statement for the binary case; the case $n\geq 3$ follows by analogy. Let $\binom{u^{1}}{v^{1}}\in R$ be min-clean, and take $I,m,n,q,$ and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in R$ as in Lemma 3.3. By induction hypothesis applied to the relation $\operatorname{pr}_{[k]\setminus I}(R)$, there exists $\binom{u}{v}\in R$ such that $u\sim_{[k]\setminus I}v$. For $\binom{u^{\prime}}{v^{\prime}}$ defined as in Lemma 3.3, we then have:
@@ 3.2. Templates with $\ell\ell$
For the proof of Lemma 3.3, we make heavy use of the fact that $pp_{q}(x,y)$ depends solely on the value of $x$ when $x\leq q$, and on the value of $y$ when $x>q$. In contrast, every operation contained in the clone $\langle\ell\ell\rangle$ is injective in all its non-dummy variables. In order to adapt to this setting, we need to ‘synchronise’ the kernels of the tuples we work with.
@@
For an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, set $R^{\prime}\coloneqq\{t\in\langle{R}\rangle_{\operatorname{lex}}\;|\;\ker t=\bigcap_{r\in R}\ker r\}.$
The relation $R^{\prime}$ inherits desirable properties from $R$:
@@
Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. Then $R^{\prime}\neq\varnothing$. If $n=2$ and $R$ is smooth, then so is $R^{\prime}$. If $\Sigma$ is a subgroup of $\textnormal{Sym}(n)$ and $R$ is $\Sigma$-invariant, then so is $R^{\prime}$.
@@
The first statement is obtained by applying $\operatorname{lex}^{[m]}$} of a sufficiently high arity $m$ to tuples in $R$ until the kernel of the resulting tuple can no more be refined. Suppose that $n=2$ and $R$ is smooth. To see that $R^{\prime}$ is smooth, pick $t\in R^{\prime}$, $f\in\langle\operatorname{lex}\rangle$ of some arity $m\in\mathbb{N}$, and $s^{1},\dots,s^{m}\in R$ such that $t=f(s^{1},\dots,s^{m}).$ We show that $t_{1}\in\operatorname{pr}_{2}(R^{\prime})$, the converse statement for $t_{2}$ is shown analogously. By smoothness of $R$, for every $i\leq m$ there exists $r^{i}\in\mathbb{Q}^{k}$ such that $\binom{r^{i}}{s^{i}_{1}}\in R$. Let $w\coloneqq f(r^{1},\dots,r^{m})$. Picking any $\binom{u}{v}\in R^{\prime}$, we now have that $\operatorname{lex}(\binom{w}{t_{1}},\binom{u}{v})\in R^{\prime}$. From $t_{1}\sim_{k}\operatorname{lex}(t_{1},v)$, the statement follows. The last statement is clear from the definition of $R^{\prime}$. ∎
For relations $R$ preserved by $\langle\ell\ell\rangle$, the existence of pseudo-loops is shown in a similar way as for relations preserved by $\langle pp\rangle$. Namely, it follows from adaptions of Lemmata 3.3 and 3.4 to this new setting (Lemmata 3.8 and 3.10 below) along with the existence of min-clean tuples (Lemmata 6.13 and 6.15, shown in Section 6). The crucial difference to the previous setting is that all tuples we work with must be contained in $R^{\prime}$. This additional prerequisite leads to the slightly weaker analogues to Lemmata 3.3 and 3.4, compensated by strengthened results showing the existence of min-clean tuples even within $R^{\prime}$.
@@
In the proof of Theorem 1.3, we apply Lemma 3.8 to a smooth weakly connected relation $S$ of pseudo-algebraic length $1$ that is contained in the edge relation of a digraph preserved by $\langle\ell\ell\rangle$. We now conclude on the existence of a pseudo-loop within $\langle{S^{\prime}}\rangle_{\ell\ell}$.
@@
Let $S$ be a binary relation on $\mathbb{Q}^{k}$. If $S$ is smooth, weakly connected and of pseudo-algebraic length $1$, then $\langle{S^{\prime}}\rangle_{\ell\ell}$ contains a pseudo-loop.
@@
Just as for templates with $pp$, a statement similar to Lemma 3.8 also applies to higher-arity relations that are cyclic.
Lemma 3.10.
Let $n\geq 2$, and let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$. If $R^{\prime}$ contains a min-clean tuple $u^{1}$, then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u\in R^{\prime}$ the tuple $u^{\prime}$ defined by
$u^{\prime}\coloneqq{\ell\ell_{q}^{[n+1]}}\left(u^{1},u^{2},\dots,u^{n},u\right)$
satisfies:
  • $I_{1}(u^{\prime}_{i})=I$ for all $i\leq n$;
  • $u^{\prime}_{i}\sim_{[k]\setminus I}u_{i}$ for all $i\leq n$.
Proof.
By Lemma 3.7, $R^{\prime}$ is cyclic, thus containing all cyclic permutations of $u^{1}$. The proof now is analogous to the one of Lemma 3.4. ∎
Under the additional assumption of $2$-transitivity, we obtain a pseudo-loop also in this setting.
@@
We proceed by induction on $k$, employing Lemma 6.15 for the existence of a min-clean tuple within $R^{\prime}$ together with Lemma 3.10. ∎
4. Proof of the main theorems
@@
Replacing $\mathbb{G}$ by the preimage $\bar{\mathbb{G}}$ of $\mathbb{G}$ under the map of the pp-interpretation in $\mathbb{A}$, we may assume that $\mathbb{G}$ is a digraph defined on a finite power of $\mathbb{Q}$ whose edge relation $E$ is preserved by every polymorphism of $\mathbb{A}$. Indeed, by definition of a pp-interpretation, the preimage $\bar{E}$ of $E$ is pp-definable in $\mathbb{A}$, and therefore preserved by $\operatorname{Pol}(\mathbb{A})$ [BN06]. A pseudo-loop in $\bar{E}$ yields a pseudo-loop in $E$ modulo $\operatorname{Aut}(\mathbb{Q})$. Moreover, $\bar{E}$ is smooth and has pseudo-algebraic length modulo $\operatorname{Aut}(\mathbb{Q})$.
By Theorem 2.3, $\mathbb{A}$ is preserved by one of $\ell\ell$, $\min$, $\operatorname{mi}$, $\operatorname{mx}$, their duals, or a constant operation. Accordingly, $\mathbb{G}$ is preserved by one of these operations, and thus by all operations in its generated clone. Clearly, every non-empty relation preserved by a constant operation contains even a loop. Moreover, $\mathbb{G}$ contains a pseudo-loop if and only if $\mathbb{G}$ does. Recalling that, for $f\in\{\ell\ell,\min,\operatorname{mi},\operatorname{mx}\}$, $\mathbb{G}$ is preserved by $\langle\operatorname{dual}(f)\rangle$ if and only if $\mathbb{G}$ is preserved by $\langle f\rangle$, this allows us to restrict ourselves to the cases where $\mathbb{G}$ is preserved by one of $\langle\ell\ell\rangle$, $\langle\min\rangle$, $\langle\operatorname{mi}\rangle$, and $\langle\operatorname{mx}\rangle$. In the first case, let $S\subseteq E$ be smooth, weakly connected, and of pseudo-algebraic length $1$ modulo $\operatorname{Aut}(\mathbb{Q})$. By Lemma 3.9, $\langle{S^{\prime}}\rangle_{\ell\ell}\subseteq E$ contains a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$. In all the other cases, we can directly apply Lemma 3.5, yielding a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$ within $E$. ∎
Similarly, putting together Lemmata 3.11 and 3.5, we obtain the corresponding pseudo-loop lemma for cyclic and $2$-transitive hypergraphs $\mathbb{T}$. It might be of interest to point out that pseudo-loop lemmata involving symmetry conditions provided by non-trivial group actions have also been established for finite hypergraphs (see [BK12, Bru23]). See 1.5
@@ 5. Pseudo-loop conditions from pseudo-loops
Following the standard method – established first by Siggers [Sig10] for finite structures, and adapted to $\omega$-categorical ones by Barto-Pinsker [BP20] – the existence of pseudo-loops in Theorems 1.3 and 1.5 implies the validity of the corresponding pseudo-loop conditions in all temporal constraint languages that are not omni-expressive. For the reader’s convenience, we provide the proof of Corollary 1.4. See 1.4
@@
Suppose that $\mathbb{G}=(\{1,\dots,n\};E_{\mathbb{G}})$ is a finite smooth digraph of algebraic length $1$, and $u\circ s(x_{i_{1}},\dots,x_{i_{m}})\approx v\circ s(x_{j_{1}},\dots,x_{j_{m}})$ is the corresponding $\mathbb{G}$-pseudo-loop condition. We show that for every $k\geq 1$ and for every $a_{1},\dots,a_{n}\in\mathbb{Q}^{k}$ there exists an $m$-ary operation $s\in\operatorname{Pol}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{Aut}(\mathbb{Q})$-orbit. A standard compactness argument (as provided, for example, in [BP20, Lemma 4.2]) using oligomorphicity of $\operatorname{Aut}(\mathbb{Q})$ then yields the ‘global’ validity in $\operatorname{Pol}(\mathbb{A})$ of the desired pseudo-identity.
To this end, take $k\geq 1$, and let $a_{1},\dots,a_{n}$ be arbitrary $k$-tuples of elements of $\mathbb{Q}$. Consider the binary relation $R\subseteq\mathbb{Q}^{k}\times\mathbb{Q}^{k}$ that consists precisely of all the tuples $(a_{i},a_{j})$ for which $(i,j)\in E_{\mathbb{G}}$, and let $E_{\mathbb{H}}$ be the smallest binary relation on $\mathbb{Q}^{k}$ that contains $R$ and is preserved by $\operatorname{Pol}(\mathbb{A})$. Since $E_{\mathbb{H}}$ is preserved by all polymorphisms of $\mathbb{A}$, it is pp-definable in $\mathbb{A}$ [BN06]. It follows that $\mathbb{H}\coloneqq(\mathbb{Q}^{k};E_{\mathbb{H}})$ is pp-interpretable in $\mathbb{A}$. Moreover, observe that $\mathbb{H}$ is a smooth digraph of algebraic length $1$. Applying Theorem 1.3, we get a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$ in $\mathbb{H}$. Unravelling definitions, this means that there exists $s\in\operatorname{Pol}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{Aut}(\mathbb{Q})$-orbit. ∎